OpenStudy (baby456):

The two functions t(x) and v(x) are defined below. t(x)=6x-1 v(x)=x^2+2 Evaluate the composition of functions v(t(4))

OpenStudy (baby456):

@ zepdrix @zzr0ck3r

OpenStudy (baby456):

@zepdrix

OpenStudy (danjs):

v(x) = x^2 + 2 so v[ t(x)] = [t(x)]^2 + 2 = [ 6x-1]^2 + 2

OpenStudy (baby456):

i dont really understand

OpenStudy (baby456):

uh where did the 6x come frome

OpenStudy (baby456):

@agent0smith could you help make it clearer

OpenStudy (agent0smith):

The 6x came from t(x)... t(x)=6x-1 v(x)=x^2+2 v[ t(x) ] = [ t(x) ]^2 + 2 = [ 6x-1 ]^2 + 2

OpenStudy (baby456):

but how about the 4

OpenStudy (baby456):

v(t(4))

OpenStudy (agent0smith):

v[ t(x) ] = [ 6x-1 ]^2 + 2 v[ t(4) ] means plug in x=4